Gabriela had to take three tests in her accounting class. She scored 71 points on
her first test and 85 points on her second tests. If her average for the three tests
is 83, what score did Gabriela get on her third test?

Found 2 solutions by solver91311, rfer:
Answer by solver91311(12126)



Multiply the average by the number of tests. Then add together the individual scores you know, subtract this sum from the product of the average and the number of tests. The result will be the score of the missing test.


John


Answer by rfer(10432)

(71+85+x)/3=83
156+x/3=83
156+x=249
x=93